# Scientific Paper Retrieval with LLM-Guided Semantic-Based Ranking

### v1

@@ Abstract
Scientific paper retrieval is essential for supporting literature discovery and research. While dense retrieval methods demonstrate effectiveness in general-purpose tasks, they often fail to capture fine-grained scientific concepts that are essential for accurate understanding of scientific queries. Recent studies also use large language models (LLMs) for query understanding; however, these methods often lack grounding in corpus-specific knowledge and may generate unreliable or unfaithful content. To overcome these limitations, we propose SemRank, an effective and efficient paper retrieval framework that combines LLM-guided query understanding with a concept-based semantic index. Each paper is indexed using multi-granular scientific concepts, including general research topics and detailed key phrases. At query time, an LLM identifies core concepts derived from the corpus to explicitly capture the query’s information need. These identified concepts enable precise semantic matching, significantly enhancing retrieval accuracy. Experiments show that SemRank consistently improves the performance of various base retrievers, surpasses strong existing LLM-based baselines, and remains highly efficient.
@@
Yunyi Zhang1, Ruozhen Yang1, Siqi Jiao1, SeongKu Kang2, Jiawei Han1 1University of Illinois Urbana-Champaign 2Korea University yzhan238@illinois.edu
@@ 1 Introduction
Recently, dense passage retrieval methods have been widely studied in various ad hoc searches Karpukhin et al. (2020); Izacard et al. (2021). These methods encode the overall semantics of queries and passages into the same vector space and measure relevance using vector similarity. Although being effective in different general domain applications, they still face challenges in scientific paper retrieval.
Specifically, general-purpose semantic representations learned by dense retrievers often fail to capture fine-grained scientific concepts that are crucial for accurately understanding and satisfying a scientific query. For example, the query “Can you point me to studies discussing methods for evaluating text generation models on various dimensions?” involves not only general topics like “natural language generation” and “automatic evaluation” that need to be inferred from the text, but also specific details like “multidimensional evaluation”. A dense retriever, however, only encodes the text in a holistic view, while it lacks the ability and controllability to focus on the scientific concepts which are the core need of the query.
@@
With the advancements of large language models (LLM) such as GPT OpenAI (2023) and Claude Anthropic (2024), recent studies also explore how to utilize LLMs in query understanding to help retrieval tasks. For example, HyDE Gao et al. (2023) uses an LLM to generate a hypothetical passage for encoding, and CSQE Lei et al. (2024) prompts an LLM to select a set of relevant sentences to expand the original query. However, these methods still rely on a pre-trained dense retriever to encode overall semantics on the document or sentence levels, lacking the ability to explicitly capture what the query is asking for. In addition, LLMs are not inherently retrieval models. They do not have the vast and dynamic knowledge in the scientific literature necessary to understand scientific queries. Therefore, when used in a zero-shot manner, LLMs cannot identify domain-specific terminology and may generate hallucinated content. Therefore, how to effectively augment LLMs with corpus-based knowledge while capturing specific query information remains a challenge.
@@ 3.2 LLM-Guided Semantic-Based Retrieval
where $\mathbf{c}$ denotes the embedding of a concept by a semantic encoder (e.g., SPECTER-v2) and $sim()$ represents the cosine similarity function. This embedding-based soft matching process identifies the most similar concept in a paper for each query’s concept, which accounts for the situation where similar concepts are expressed slightly differently (e.g., "hallucination" and "hallucinated content"). Then, we combine this semantic-based score with the base retriever’s score by z-score normalization (denoted by $z(\cdot)$), based on which we can re-rank the papers to get a new ranked list.
@@ 4.1 Experiment Setup
We use a wide range of base retrievers in our experiments. We include a sparse retriever BM25, an unsupervised dense retriever SPECTER-v2 Singh et al. (2022), and the Hybrid of these two. We also include two instruction-tuned dense retrievers, E5-large-v2 Wang et al. (2022) and GritLM-7B Muennighoff et al. (2025).
@@
For building semantic index, we train our text classifier using MAPLE Zhang et al. (2023b) which contains topic labels from Microsoft Academic Graph Sinha et al. (2015). We initialize it with SPECTER-v2-base and fine-tune it with learning rate at 5e-5 for 10 epochs. The balancing factor $\alpha$ is set to 1e-2. We also use SPECTER-v2-base for concept encoding. During retrieval, the number of candidate topics and key phrases is $k=50$. We use GPT-4.1-mini as the LLM for SemRank and all LLM-based baselines for fair comparison. The experiments are run on one NVIDIA RTX A6000 when a GPU is needed.  [^4] Code is available upon request and will later be public.
@@ 4.4 Efficiency Analysis
We show the efficiency of SemRank by comparing it with other LLM-based baselines. Specifically, we report the following factors of each method: the number of base retriever calls per query (# RET), the number of LLM calls per query (# LLM), the average number of tokens generated by the LLM per query (LLM Output Len), and the average running time per query (Running Time). As stated in Sect. 4.1, we use the same LLM checkpoint for all baselines for a fair comparison. Table 5 shows the detailed comparison results on LitSearch. We can clearly see that SemRank takes the least inference time among the compared methods, with 1.5$\times$ faster than the second fast method. Not only does SemRank call the retriever and LLM only once, it also expects minimal number of tokens responded by the LLM because of the concept-based structured input and output format.
@@ 5 Related Works
In academic domains, specialized models like SciBERT Beltagy et al. (2019) pre-trained on scientific texts laid the groundwork. Recent innovations also focus on knowledge distillation from cross-encoder rankers  Huang and Chen (2024); Tao et al. (2024); Zhang et al. (2023a). Despite the progress and varied strategies for improving dense retrieval, these methods inherently rely on representing entire documents or queries with single dense embeddings, restraining them to capture fine-grained details crucial for accurately interpreting complex scientific queries. In contrast, our approach explicitly models the multi-granular scientific concepts within both queries and documents.
@@ 6 Conclustion
While SemRank shows strong performance and efficiency in scientific paper retrieval, it also has several limitations. First, our current studies limit to scientific paper retrieval dataset with only title and abstract, while retrieving full scientific papers could be more challenging due to the difficulties of effectively understanding long structured text. Second, SemRank only consider scientific concepts as a set, while not considering their internal relationship which could bring more insights to paper and query understanding. Third, although our use of LLMs is efficient, the reliance on prompting still introduces sensitivity to prompt design and model behavior, which may require tuning for different domains. Finally, we only focus on English scientific paper retrieval in the work, while it remains a challenge on multilingual or multi-modal (e.g., figures, tables) paper retrieval.
@@ Acknowledgments
Research was supported in part by US DARPA INCAS Program No. HR0011-21-C0165 and BRIES Program No. HR0011-24-3-0325, National Science Foundation IIS-19-56151, the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No. 2019897, and the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under Award No. 2118329.

### v2

@@ Abstract
Scientific paper retrieval is essential for supporting literature discovery and research. While dense retrieval methods demonstrate effectiveness in general-purpose tasks, they often fail to capture fine-grained scientific concepts that are essential for accurate understanding of scientific queries. Recent studies also use large language models (LLMs) for query understanding; however, these methods often lack grounding in corpus-specific knowledge and may generate unreliable or unfaithful content. To overcome these limitations, we propose SemRank, an effective and efficient paper retrieval framework that combines LLM-guided query understanding with a concept-based semantic index. Each paper is indexed offline using multi-granular scientific concepts, including general research topics and detailed key phrases. At query time, an LLM identifies core concepts derived from the corpus to explicitly capture the query’s information need. These identified concepts enable precise semantic matching, significantly enhancing retrieval accuracy. Experiments show that SemRank consistently improves the performance of various base retrievers, surpasses strong LLM-based baselines, and remains highly efficient. [^1] Code can be found at: https://github.com/yzhan238/SemRank.
@@
Yunyi Zhang1, Ruozhen Yang1, Siqi Jiao1, SeongKu Kang2, Jiawei Han1 1University of Illinois Urbana-Champaign 2Korea University {yzhan238,ruozhen2,sjiao2,hanj}@illinois.edu seongkukang@korea.ac.kr
@@ 1 Introduction
Recently, dense passage retrieval methods have been widely studied in various ad-hoc searches Karpukhin et al. (2020); Izacard et al. (2021). These methods encode the overall semantics of queries and passages into the same vector space and measure relevance using vector similarity. Although being effective in different general domain applications, they still face challenges in scientific paper retrieval.
Specifically, general-purpose semantic representations learned by dense retrievers often fail to capture fine-grained scientific concepts that are crucial for accurately understanding and satisfying a scientific query Chen et al. (2022); Shavarani and Sarkar (2025). For example, the query “Can you point me to studies discussing methods for evaluating text generation models on various dimensions?” involves not only general topics like “natural language generation” and “automatic evaluation” that need to be inferred from the text, but also specific details like “multidimensional evaluation”. A dense retriever, however, only encodes the text in a holistic view, while it lacks the ability and controllability to focus on the scientific concepts which are the core need of the query.
@@
With the advancements of large language models (LLM) such as GPT OpenAI (2023) and Claude Anthropic (2024), recent studies also explore how to utilize LLMs in query understanding to help retrieval tasks. For example, HyDE Gao et al. (2023) uses an LLM to generate a hypothetical passage for encoding, and CSQE Lei et al. (2024) prompts an LLM to select a set of relevant sentences to expand the original query. However, these methods still rely on a pre-trained dense retriever to encode overall semantics on the document or sentence levels, lacking the ability to explicitly capture what the query is asking for. In addition, LLMs are not inherently retrieval models. They do not have the vast and dynamic knowledge in the scientific literature necessary to understand scientific queries. Therefore, when used in a zero-shot manner, LLMs cannot identify domain-specific terminology and may generate hallucinated content. Thus, how to effectively augment LLMs with corpus-based knowledge while capturing specific query information remains a challenge.
@@ 3.1 Semantic Index Construction
By first predicting candidate topic labels with a text classifier, not only can we resolve the issue of LLMs’ poor performance on a large structured label space U et al. (2023); Zhang et al. (2025), but also reduce the cost of prompting LLMs by 98%. A more detailed cost analysis is in Section 4.4. Also note that we only need to train one topic classifier for a domain and use it for all corpora in the same domain (e.g., Computer Science).
@@ 3.2 LLM-Guided Semantic-Based Retrieval
where $\mathbf{c}$ denotes the embedding of a concept by a semantic encoder (e.g., SPECTER-v2) and $sim()$ represents the cosine similarity function. This embedding-based soft matching process identifies the most similar concept in a paper for each query’s concept, which accounts for the situation where similar concepts are expressed slightly differently (e.g., “hallucination” and “hallucinated content”). Then, we combine this semantic-based score with the base retriever’s score by z-score normalization (denoted by $z(\cdot)$), based on which we can re-rank the papers to get a new ranked list.
@@ 4.1 Experiment Setup
We use a wide range of base retrievers in our experiments. We include a sparse retriever BM25, an unsupervised dense retriever SPECTER-v2 Singh et al. (2022), and the Hybrid of these two. We also include two instruction-tuned dense retrievers, E5-large-v2 Wang et al. (2022) and GritLM-7B Muennighoff et al. (2025) [^5] We also experiment with other retrievers such as NV-Embed-v2 Lee et al. (2025). GritLM achieves the strongest performance on scientific paper retrieval, potentially because it specifically includes scientific corpora in its training data..
@@
  • Boudin et al. Boudin et al. (2020) uses a seq2seq keyphrase generation model to enrich the corpus indexing.
@@
For building semantic index, we train our text classifier using MAPLE Zhang et al. (2023b) which contains topic labels from Microsoft Academic Graph Sinha et al. (2015). We initialize it with SPECTER-v2-base and fine-tune it with learning rate at 5e-5 for 10 epochs. The balancing factor $\alpha$ is set to 1e-2. We also use SPECTER-v2-base for concept encoding. During retrieval, the number of candidate topics and key phrases is $k=50$. We use GPT-4.1-mini as the LLM for SemRank and all LLM-based baselines for fair comparison. The experiments are run on one NVIDIA RTX A6000 when a GPU is needed.
@@ 4.4 Efficiency Analysis
[…]
Additionally, the offline indexing part of SemRank is efficient as well. For semantic index construction on CSFCube, to get the candidate topics, the text classifier inference time is 52 seconds, and to prompt LLM for topic and keyphrase selection, it takes 84 seconds and \$1.74 dollars. In comparison, directly prompting LLMs to build such semantic index on CSFCube will take approximately 42 minutes and \$85 dollars.
@@ 5 Related Works
In academic domains, specialized models like SciBERT Beltagy et al. (2019) pre-trained on scientific texts laid the groundwork. Parisot and Zavrel (2022) proposes a multi-objective approach that uses general-domain document relevance and scientific domain citation network and self-supervised data. Mandikal and Mooney (2024) presents a hybrid approach that combines sparse and dense retrievers with a weighting parameter. MixGR Cai et al. (2024) matches queries and documents additionally at subquery and proposition levels and merge them with rank fusion. Recent innovations also focus on knowledge distillation from cross-encoder rankers  Huang and Chen (2024); Tao et al. (2024); Zhang et al. (2023a). Despite the progress and varied strategies for improving dense retrieval, these methods inherently rely on representing entire documents or queries with single dense embeddings, restraining them to capture fine-grained details crucial for accurately interpreting complex scientific queries. In contrast, our approach explicitly models the multi-granular scientific concepts within both queries and documents.
@@ 6 Conclusion
While SemRank shows strong performance and efficiency in scientific paper retrieval, it also has several limitations. First, our current studies limit to scientific paper retrieval dataset with only title and abstract, while retrieving full scientific papers could be more challenging due to the difficulties of effectively understanding long structured text. Second, SemRank only considers scientific concepts as a set, while not considering their internal relationship which could bring more insights to paper and query understanding. Third, although our use of LLMs is efficient, the reliance on prompting still introduces sensitivity to prompt design and model behavior, which may require tuning for different domains. Fourth, our experiments are done mainly on Computer Science domain corpora, because there is limited high-quality retrieval dataset available from other disciplines. We would like to argue that there is still a big gap in constructing paper retrieval benchmarks for different disciplines, which could be a research opportunity for future studies. Finally, we only focus on English scientific paper retrieval in the work, while it remains a challenge on multilingual or multi-modal (e.g., figures, tables) paper retrieval.
@@ Acknowledgments
Research was supported in part by National Science Foundation IIS-19-56151, NSF IIS 25-37827, the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No. 2019897, and the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under BRIES Program No. HR0011-24-3-0325. The research has used the Delta/DeltaAI advanced computing and data resource, supported in part by the University of Illinois Urbana-Champaign and through allocation #250851 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants OAC 2320345, #2138259, #2138286, #2138307, #2137603, and #2138296. Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views, either expressed or implied, of DARPA or the U.S. Government.